# D-Band 2D MIMO FMCW Radar System Design for Indoor Wireless Sensing

### v3

@@ I-E Angular Resolution
For example, by doubling the number of TX or RX elements, the SNR of the MIMO radar can be improved by 3 dB.
II DOA Estimation Using 2D Radar Array
II-A MUSIC and MVDR Algorithms
@@
The covariance matrix and the array matrix are used to construct a spatial spectrum function $P(\theta,\phi)$. In the MUSIC algorithm, $P(\theta,\phi)$ is related to the signal and noise subspace of the covariance matrix [25], while in the MVDR algorithm it is derived from the inverse of the covariance matrix [26]. Finally, a search algorithm is used to find peaks of the spatial spectrum function as DOA of the source signals.
II-B Scaling of SNR and Output Power with Distance
In free space, the SNR of the received signal is derived as
where $P_{TX}$ is the TX output power, $G_{TX}$ and $G_{RX}$ are gain of the TX and RX antennas, $\lambda$ is the signal wavelength, $\sigma$ is the radar cross section (RCS) of the target, $T_{meas}$ is the radar measurement time, $k=\rm{1.38\times 10^{-23}\,J/K}$, $T$ is the absolute temperature, $F$ is the RX noise factor, $\rm{NF_{RX}=10\log_{10}(F)}$, and $d$ is the target distance.
In classical DOA estimation problems, the SNR is considered as an independent parameter, and, as a result, it is independent of the target distance. However, this requires scaling of the TX output power with the distance as $P_{\rm{TX}}\propto d^{4}$ [see (15)]. This can lead to an impractical TX output power requirement at long distances. On the other hand, if we assume a fixed TX output power, the RX SNR will decrease with the distance as ${\rm{SNR}}\propto d^{-4}$. A compromise between the two extremes is to assume that both the TX output power and the RX SNR are scaled with the distance.
We propose a scaling approach for the RX SNR and TX output power with the distance as
${\rm{SNR}}(d)={\rm{SNR}}(d_{0})\left(\frac{d_{0}}{d}\right)^{p}$ (16)
$P_{\rm{TX}}(d)=P_{\rm TX}(d_{0})\left(\frac{d_{0}}{d}\right)^{p-4},$ (17)
where $p$ is the scaling exponent that should be selected in the range $0\leq p\leq 4$ and $d_{0}$ is the reference distance. The TX output power is set to achieve a certain RX SNR at a reference distance $d_{0}$, and it is scaled according to (17) at other distances. The classic constant SNR scenario is achieved for $p=0$, while the case that the TX output power is constant corresponds to $p=4$. Different conditions for the scaling of the SNR and TX output power with distance using (16) and (17) are presented in Fig. 4. We select $p=2$ as a compromise between the slope of the SNR drop with distance and the increase in the required TX power.
Figure 4: Scaling of the RX SNR and the TX output power with distance. In the fixed SNR scenario (blue), the TX power should be increased with the slope of $+40\log(d/d_{0})$, while in the fixed TX power scenario (red), the SNR is reduced with the slope of $-40\log(d/d_{0})$. A compromise can be achieved by selecting a scaling exponent in the range $40<p<$ 4.
II-C DOA Estimation Results
@@
The MUSIC and MVDR DOA estimation algorithms are evaluated for different 2D radar array architectures. The reference distance is assumed as $d_{0}={\rm{1\,m}}$. Root-mean-square error (RMSE) is used as the accuracy metric to evaluate the algorithms
${\rm{RMSE}}=\sqrt{\frac{1}{KM}\sum_{j=1}^{M}\sum_{i=1}^{K}\left((\hat{\phi_{i}}(j)-\phi_{i})^{2}+(\hat{\theta_{i}}(j)-\theta_{i})^{2}\right)},$ (18)
@@
where $(\phi_{i},\theta_{i})$ is the actual DOA of the i-th target, $(\hat{\phi_{i}}(j),\hat{\theta_{i}}(j))$ is the estimated DOA of the i-th target in the j-th sample of the Monte Carlo runs, $K$ is the number of targets, and $M$ is the number of the Monte Carlo runs. We model the noise present in the received signal as a random Gaussian variable, $\sim\mathcal{N}(0,\sigma^{2})$, with zero mean and a variance determined based on the SNR. We evaluated the simulation results for different numbers of runs and selected $M=100$ to achieve high accuracy with practical simulation times. In this work, we assume that there are three targets, $K=3$, with equal signal amplitudes.
Figure 5: Simulation times of the DOA estimation for different algorithms and array sizes.
II-C1 64-Element Array
@@
For a 64-element array realized as an $8\times 8$ square array, the RMSE in terms of SNR and distance is shown in Fig. 6. The SNR and the distance are, respectively, varied in the range of 0–30 dB and 1–10 m, as typical values in indoor wireless sensing scenarios. The MUSIC algorithm outperforms the MVDR algorithm in the all SNR levels and distances, especially at long-distance and low-SNR conditions. Fig. 6(a) indicates that at short distances, for example 1 m, the RMSE lower than $1^{\circ}$ can be maintained even at very low SNR levels. This is the result of an improved SNR of the MIMO array that for the 64-element array is derived using (11) as $\rm 10\log_{10}(64)\approx 18\,dB$. As shown in Fig. 6(b), for the SNR of 20 dB, RMSE lower than $0.6^{\circ}$ can be achieved across the 1–10 m distances using the MUSIC algorithm.
A 64-element array can be realized as other structures, for example, $16\times 4$, $32\times 2$, as a higher resolution in the azimuth angle compared to the elevation angle is usually required. Simulation results of these alternative array structures are not shown due to limited space. The DOA estimation results for the $16\times 4$ array indicate RMSE of about $1.2-1.6^{\circ}$ for SNR of 0–30 dB at 1 m distance. Compared to the RMSE of the $8\times 8$ array shown in Fig. 6(a), the RMSE of the $16\times 4$ array is higher by 2x–3x. This can be due to reduced number of array elements in the vertical direction which leads to a lower elevation resolution.
Simulation times of the DOA estimation are shown in Fig. 5. Simulations are performed using an Intel Core i5-7300HQ 2.50 GHz CPU with 8 GB RAM. The simulation time comprises the time required to complete the simulations on an SNR of 0–30 dB with a step of 5 dB, a distance of 1–10 m with a step of 1 m, and 100 Monte Carlo runs. For the $8\times 8$ array, the MUSIC algorithm provides faster simulations and higher accuracy compared to the MVDR algorithm.
@@
Figure 6: RMSE for the MUSIC and MVDR algorithms with $8\times 8$ radar array. MUSIC outperforms MVSR especially in the low-SNR conditions.
II-C2 16-Element Array
@@
For a 16-element array realized as a $4\times 4$ structure, the RMSE in terms of the SNR and distance is shown in Fig. 7. Interestingly, the MUSIC and MVDR algorithms provide very close accuracy across the ranges of SNR and distance considered in the simulations. This is in contrary with the $8\times 8$ array in which the MUSIC led to higher accuracy compared to the MVDR. The SNR improvement by the MIMO array is $\rm 10\log_{10}(16)\approx 12\,dB$. Fig. 7(b) indicates that with the 20 dB SNR, the RMSE lower than $2^{\circ}$ can be achieved in the 1–10 m range. Moreover, Fig. 5 shows that the simulation times for the two algorithms are very close.
@@
Figure 7: RMSE for the MUSIC and MVDR algorithms with $4\times 4$ radar array. MUSIC and MVDR provide very close results.
II-C3 4-Element Array
A 4-element array can only provide a coarse DOA estimation, but its simpler hardware and lower power consumption can be very useful for new applications (for example, smart home appliances and health monitoring devices). The SNR boost by the MIMO structure is $\rm 10\log_{10}(4)\approx 6\,dB$ for the 4-element array. The RMSE for a $2\times 2$ array versus SNR and distance is shown in Fig. 8. In this case, the MVDR algorithm significantly outperforms the MUSIC algorithm. This is an important result indicating that the MVDR algorithm is more accurate for small radar arrays. Fig. 8(b) shows that with 20 dB SNR, RMSE lower than $6^{\circ}$ can be achieved in 1–10 m using the MVDR algorithm. A higher SNR of 30 dB can reduce the RMSE to $0.4^{\circ}$ at 1 m and $1.3^{\circ}$ at 5 m [Fig. 8(a)]. Furthermore, Fig. 5 indicates that MVDR also features a simulation time 0.8x shorter than that of the MUSIC algorithm.
Figure 8: RMSE for the MUSIC and MVDR algorithms with $2\times 2$ radar array. MVDR outperforms MUSIC across all SNR and distance ranges.
II-C4 Accuracy Versus the Number of Array Elements
In Fig. 9, the RMSE is shown versus the number of array elements. The minimum RMSE achieved using the MUSIC and MVDR algorithms in each condition is presented here. This can be used to determine the number of elements of the array to achieve a desired RMSE for a specific SNR and distance.

### v4

@@ II-B Adaptive Radar Resolution
For example, by doubling the number of TX or RX elements, the SNR of the MIMO radar is improved by 3 dB. The accuracy of the DOA estimation is dependent on the number of array elements, the SNR of the received signal, and the algorithm. Therefore, an adaptive radar resolution can be realized by varying the number of array elements.
Figure 1: Architecture of the 2D MIMO radar comprising $N_{x}\times N_{y}$ virtual arrays. The effective array size can be adaptively reconfigured by disabling certain TX and RX elements to improve the energy efficiency.
II-C Scaling of SNR and Output Power with Distance
In free space, using the fundamental radar range equation, the SNR of the received signal can be derived as
where $P_{TX}$ is the TX output power, $G_{TX}$ and $G_{RX}$ are gain of the TX and RX antennas, $\lambda$ is the signal wavelength, $\sigma$ is the radar cross section (RCS) of the target, $T_{meas}$ is the radar measurement time, $k=\rm{1.38\times 10^{-23}\,J/K}$, $T$ is the absolute temperature, $F$ is the RX noise factor, and $d$ is the target distance.
In classic DOA estimation approaches, the SNR is considered as a parameter independent of the target distance. However, this approach requires scaling of the TX output power with the distance as $P_{\rm{TX}}\propto d^{4}$ [see (7)]. This can lead to an impractical TX output power requirement at long distances. On the other hand, if we assume a fixed TX output power, the RX SNR will decrease with the distance as ${\rm{SNR}}\propto d^{-4}$. A compromise between the two extremes is to assume that both the TX output power and the RX SNR are scaled with the distance.
We propose a scaling approach for the RX SNR and the TX output power with the distance as
${\rm{SNR}}(d)={\rm{SNR}}(d_{0})\left(\frac{d_{0}}{d}\right)^{p}$ (8)
$P_{\rm{TX}}(d)=P_{\rm TX}(d_{0})\left(\frac{d_{0}}{d}\right)^{p-4},$ (9)
where $p$ is the scaling exponent that should be selected in the range $0<p<4$ and $d_{0}$ is the reference distance. The TX output power is set to achieve a certain RX SNR at a reference distance $d_{0}$, and it is scaled according to (9) at other distances. The classic constant SNR scenario is achieved for $p=0$, while the case that the TX output power is constant corresponds to $p=4$. In Fig. 2, the profiles of RX SNR and TX output power versus distance are shown for values of $p$ using (8) and (9). We select $p=2$ for a compromise between the slopes of the SNR and the required TX output power.
Figure 2: Scaling of the RX SNR and the TX output power with distance. In the fixed SNR scenario (blue), the TX power should be increased with the slope of $40\log(d/d_{0})$, while in the fixed TX power scenario (red), the SNR is reduced with the slope of $40\log(d/d_{0})$. A compromise can be achieved by selecting a scaling exponent in the range $0<p<4$.
II-D Adaptive MIMO Waveforms
In the MIMO radar, orthogonal waveforms are transmitted to simplify the separation of received signals [27]. This can be achieved by applying a multiplexing scheme to modulate the chirp sequence signal $S_{\rm{TX}}(t)$ into the TX signals, $S_{\rm{TX,k}}(t)$ for $k=1,...,N_{\rm{TX}}$, as shown in Fig. 3. Time-division multiplexing (TDM) is the simplest scheme for MIMO radar where only one TX transmits at a time slot
$x_{k,{\rm TDM}}(t)=p\left(\frac{t-\Delta t_{k}}{T}\right),$ (10)
where $p(t)$ is the pulse function. The time offset is selected as $\Delta t_{k}=(k-1)T/N_{TX}$, which should be adaptively varied based on the number of active TX elements. However, the TDM scheme degrades the overall transmitted power, which is crucial for mm-wave bands.
In the Doppler-division multiplexing (DDM) scheme, a phase code is applied with a certain pattern to realize the TX signals
$x_{k,{\rm DDM}}(t)=e^{-j\Delta\Phi_{k}(t)},$ (11)
where the phase offset can have various forms. A simple approach is to select it as $\Delta\Phi_{k}(t)=(k-1)2\pi/N_{TX}$, which should be adaptively varied based on the number of active TX elements.
In the orthogonal frequency-division multiplexing (OFDM) scheme, the TX signal is modulated by different carrier frequencies, which are usually shifted by an offset frequency
$x_{k,{\rm OFDM}}(t)=e^{-j[2\pi(k-1)\Delta f]t}.$ (12)
The offset frequency should be adjusted based on the number of active TX elements if a certain bandwidth is considered for the OFDM scheme, $\Delta f=B_{\rm OFDM}/N_{TX}$.
In the DDM and OFDM MIMO schemes, all of the radar transmitters operate simultaneously and the total output power is enhanced. Therefore, we assume that the MIMO radar is realized using DDM or OFDM schemes with adaptive waveforms varied by the number of active TX elements.
Figure 3: Conceptual orthogonal waveform generation for MIMO FMCW radar. The chirp signal $S_{\rm{TX}}(t)$ is modulated by the sequence $x_{k}(t)$ to realize the transmit signals $S_{\rm TX,k}(t)$ for $k=1,...,N_{\rm{TX}}$. The MIMO waveforms are adaptively varied as the number of TX elements change.
III MIMO Radar Front-End
III-A Front-End Architecture
We consider the MIMO radar front-end shown in Fig. 4, which comprises two uniform linear arrays of the RX and TX elements. This is equivalent to a 2D virtual array of $N_{\rm TX}N_{\rm RX}$ elements. This architecture enables a symmetric LO distribution to the TX and RX elements. This leads to correlated phase noise of the TX and RX signals, which in a FMCW radar suppresses the phase noise in the received IF signal [6]. The LO distribution network is a critical component of a MIMO radar, especially in the D-band, and can dominate the power consumption of the overall radar system. Therefore, it is imperative to minimize the associated losses and power consumption. We use this architecture to estimate the losses and power consumption of the LO distribution network, the required TX output power, and the power consumption of the MIMO radar front-end.
The D-band antenna array can be implemented using the antenna-in-package (AiP) technology and directly flip-chip attached to the chip die [29]. This allows the integration of large MIMO radars with high radiation efficiencies (e.g., over 80%).
Figure 4: Front-end architecture of the 2D MIMO radar realized using two linear arrays of the TX and RX elements. The LO signal distribution network comprises the losses of the LO signal paths $L_{P}$, the losses of the power dividers $L_{D}$, and the LO amplifiers with power gain of $G_{\rm LO,TX}$ and $G_{\rm LO,RX}$.
III-B LO Signal Distribution
The LO generation comprises a phase-locked loop (PLL) and frequency multiplier(s) with low spurs, wide bandwidth, good chirp linearity, and high DC-to-RF efficiency. The reference signal can be generated by an external PLL. The LO distribution network, as shown in Fig. 4, comprises the routing paths to the TX and RX elements, and the LO amplifiers to compensate for the losses and maintain the signal integrity.
Frequency multipliers can be used at different locations in the MIMO radar to upconvert the low-frequency reference signal to the D-band LO signals. We can envision two extreme cases for the deployment of frequency multipliers. In the first approach, the reference signal is unconverted using a single high-performance frequency multiplier (usually comprising multiple stages) to the final RF signals and is then distributed between the TX and RX systems. The drawback is the higher loss of the LO distribution network and the higher power consumption of the LO amplifiers. In the second approach, a low-frequency reference signal is distributed between the TX and RX paths, and is upconverted by dedicated frequency multipliers in the all TX and RX systems. This can reduce the losses of the LO distribution network and the power consumption of the LO amplifiers. However, this needs a total of $N_{TX}+N_{RX}$ frequency multipliers which can lead to high power consumption for large MIMO radars. The optimal architecture should be between these two extremes, but it depends on many factors on the circuit and process level. In this work, we use the first approach to perform system simulations.
In Fig. 4, the output power of the reference signal after being upconverted by the frequency multiplier is $P_{\rm REF}$, and after power division between the routing paths it reaches to $P_{\rm LO,TX}$ in the TX and $P_{\rm LO,RX}$ in the RX systems. Both of these LO power levels should be within certain ranges for their efficient operation. In the TX, the LO power should be high enough to push the PA into saturation and achieve high efficiency, as shown in Fig. 5(a). However, an excessively high LO power can degrade gain, power-added efficiency (PAE), defined as $\rm{PAE=(P_{out}-P_{in})/P_{DC}}$, and the linearity of the PA. In the RX, the LO power should be high enough to enable switching of the transistors operating as the mixer, as shown in Fig. 5(b). On the other hand, for excessively large LO power levels, the conversion gain of the mixer, defined as $\rm{G_{c}=P_{out,IF}/P_{in,RF}}$, will degrade due to the operation of the transistor pair in the triode mode during most of the signal cycle [10]. Therefore, the LO power delivered to TX and RX should be maintained within optimal power ranges,
$P_{\rm LO,min}\leq P_{\rm LO}\leq P_{\rm LO,max},$ (13)
where the lower and upper limits of the LO power are dependent on the properties of the circuit and fabrication process.
The reference signal is highly attenuated by the divisions between the routing paths and the losses of on-chip transmissions lines in the D-band. Therefore, the LO amplifiers are distributed along the routing paths to prevent the need for a single high-gain LO amplifier after the reference signal generator. We propose to use one LO amplifier in each path with a power gain of $G_{\rm LO,TX}$ in the TX paths and $G_{\rm LO,RX}$ in the RX paths (see Fig. 4). The loss of the LO paths is dependent on their physical length, layout structure, and properties of the chip fabrication process. We assume that the length of all branches is $\Delta l$ to simplify derivations. This length is dependent on the routing layout, but it can be fairly approximated by a fraction of the signal wavelength, e.g.,, $\Delta l\approx\lambda/2$. The loss per unit length of the LO distribution paths can be estimated using the loss properties of transmission lines in the CMOS process (typically 1–3 dB/mm in the D-band). The loss of each path with the length of $\Delta l$ is derived as $L_{P}$.
The LO power delivered to each of the TX and RX systems can be derived using Fig. 4 as
${P_{\rm LO,TX}}=\frac{L_{D}(G_{\rm LO,TX}L_{P})^{\log_{2}{N_{TX}}}}{N_{TX}}P_{REF}$ (14)
${P_{\rm LO,RX}}=\frac{L_{D}(G_{\rm LO,RX}L_{P})^{\log_{2}{N_{RX}}}}{N_{RX}}P_{REF},$ (15)
The gain of the LO amplifiers can be adjusted to compensate for the losses or amplify the reference signal if required.
The total number of the LO amplifiers in the TX path can be calculated using Fig. 4 as
which using $\sum_{k=0}^{n}a^{k}=(a^{n+1}-1)/(a-1)$ is derived as
$N_{\rm LO,TX}=2N_{TX}-1.$ (19)
Similarly, in the RX path, the number of the LO amplifiers can be derived as
$N_{\rm LO,RX}=2N_{RX}-1.$ (20)
These results will be used to estimate the power consumption of the LO distribution network in Section III-D.
Figure 5: Structure of typical $m$-wave circuits used in the radar: (a) power amplifier circuit, and its efficiency $\eta$ and power-added efficiency (PAE) versus input power, (b) mixer circuit and its conversion gain (CG) versus LO power, (c) amplifier circuit with power gain boosting, and the maximum power gain $G_{\rm max}$, unilateral power gain $U$, and the power gain efficiency (PGE), $\alpha=G_{P}/{P_{DC}}$}, for an amplifier versus power consumption.
III-C Power Consumption of Radar Front-End
The power consumption of the MIMO radar is dependent on the number of array elements, system architecture of the radar TX and RX, LO distribution network, circuit structures, and fabrication process. In the D-band, the power consumption of the RX circuits is comparable with that of the TX circuits, and the LO distribution network can have a significant contribution which dominates the total power consumption. The power consumption of the radar front-end can be estimated as
$P_{\rm{DC}}\approx P_{\rm{DC,TX}}+P_{\rm{DC,RX}}+P_{\rm{DC,LO}}.$ (21)
III-C1 TX Power Consumption
The required TX output power for each TX element of the array can be derived using (7), (8), and (9) as
This indicates that for $p<4$, the required TX output power increases with the distance (see Fig. 2). Conventional DOA estimation approaches assume that the SNR is independent of the target distance, that is $p=0$. However, (22) indicates that this requires the TX power to be increased proportional with $d^{4}$. This can result in high power consumption in long indoor distances. In our proposed approach, we scale both the SNR and the TX output power with distance by selecting $p=2$. This can significantly reduce the required TX output power and the power consumption of the MIMO radar.
$P_{\rm{DC,TX}}(d)\approx\frac{N_{\rm{TX}}P_{\rm{TX}}(d)}{\eta_{\rm{TX}}}.$ (23)
In the MIMO radar architecture considered in this work, the TX efficiency can be approximated by the PA efficiency, which is in the order of 5–10% for silicon-based PAs in the D-band [6, 8, 10, 11, 13, 30]. It is assumed that the TX circuits are reconfigured as the TX output power is scaled with the distance so that the efficiency remains close to the PA peak efficiency. For the FMCW radars where a saturated PA can be used, this is less challenging than the communication systems using the amplitude-modulated signals. As a result, the power consumption of the TX system also scales with the distance.
III-C2 RX Power Consumption
The power consumption of these circuits is dependent on many factors on the circuit and process level and, as a result, cannot be accurately predicted before the circuits are designed. We can assume a fixed power consumption for each RX and scale it with the number of RX elements in the array. To find a practical estimate, we investigated the power consumption of the TX, RX, and LO parts in the D-band radars presented in the literature and noted that the power consumption of each RX element is usually on the order of 0.5x–1x the power consumption of each TX element [6, 9, 10, 12]. Therefore, we estimate the power consumption of an RX element using that of an TX element as $P_{\rm{DC,RX}}/N_{\rm{RX}}\approx kP_{\rm{DC,TX}}/N_{\rm{TX}}$ with $k\approx 0.75$. We should note that the RX power consumption is not scaled with the TX output power (unlike the TX). Therefore, $P_{\rm{DC,TX}}$ should be evaluated using (23) at a fixed target distance, such as the reference distance $d_{0}$.
III-C3 LO Power Consumption
The LO distribution network comprises the LO amplifiers and frequency multiplier(s). The LO amplifiers are designed to achieve the maximum power gain, which is different with the design criteria of the PA and LNA in the TX and RX elements. We can define an efficiency metric namely the power gain efficiency for an LO amplifier as the ratio of the power gain to the power consumption
$\alpha=\frac{G_{P}}{P_{DC}}.$ (25)
In mm-wave bands above 100 GHz, the gain of transistors is very low and usually an embedding passive network is used to boost it [31, 32]. The maximum power gain of the transistor is given by
$G_{\rm max}=(\sqrt{U}+\sqrt{U-1})^{2},$ (26)
where $U$ is the unilateral power gain (Mason’s invariant), dependent on the size and bias of the transistor and the operating frequency [31]. By increasing the bias current, $U$ initially increases but after a certain current level it is saturated [see Fig. 5(c)]. Therefore, we can assume that the transistor is biased at an optimum current with an associated unilateral power gain of $U_{opt}$. Using (25) and (26), we can conclude that this optimal current leads to the maximum $\alpha$ [see Fig. 5(c)]. Usually, multiple amplifier stages should be cascaded to achieve the desired power gain.
$P_{\rm DC,LO}=P_{M}+\frac{(2N_{TX}-1)G_{\rm LO,TX}}{\alpha_{\rm LO,TX}}+\frac{(2N_{RX}-1)G_{\rm LO,RX}}{\alpha_{\rm LO,RX}},$ (27)
where $P_{M}$ is the power consumption of the frequency multiplier(s). This equation along with (16) and (17) indicate that the power consumption of the LO network is dependent on the number of TX and RX elements, the power gain from the reference to the LO signals, that is $P_{\rm TX,LO}/P_{\rm REF}$ and $P_{\rm RX,LO}/P_{\rm REF}$, the losses of the passive components, $L_{P}$ and $L_{D}$, and the power gain efficiency of the transistors, $\alpha_{\rm LO,TX}$ and $\alpha_{\rm LO,RX}$. As we have considered a front-end architecture with a single frequency multiplier, $P_{M}$ is not scaled with the number of TX and RX elements.
III-C4 System Simulations
The power consumption of the TX, RX, and LO distribution network versus the number of MIMO radar elements $N_{\rm TX}N_{\rm RX}$ is shown in Fig. 6. The following assumptions are applied. The TX output power and efficiency are 10 dBm and 10%, the target distance is 1 m, and the power consumption of the frequency multiplier is 50 mW. The LO distribution network provides 10 dB power gain from the reference signal to the LO signals, i.e., $P_{\rm LO,TX}/P_{\rm REF}=P_{\rm LO,RX}/P_{\rm REF}=10$ in (16) and (17). The losses of the LO paths and the power dividers are $L_{P}=2\,{\rm dB}$ and $L_{D}={\rm 1\,dB}$. The power consumption of the LO distribution network is shown for two values of the power gain efficiency.
Fig. 6 indicates that the power consumption of the LO distribution network is dominant when the number of array elements is small. In this condition, each LO amplifier should provide a high gain to meet the required gain from the reference signal to the LO signals. By increasing the number of array elements, two competing effects should be considered. First, the required gain of each LO amplifier is lower, as more LO amplifiers are cascaded and higher gain is achieved. This effect tends to reduce the power consumption of the LO amplifiers. Second, as the number of LO amplifiers increases with the number of array elements, (19) and (20), this effect tends to increase the power consumption. These two competing trends initially reduce the LO power consumption, leading to a flat area shown in Fig. 6, and then increase the LO power consumption as the second effect dominates.
The power consumption of the TX and RX systems increases with the number of array elements (assuming $N_{\rm TX}=N_{\rm RX}$). For large number of array elements, the power consumption is dominated by the TX power consumption when the power gain efficiency is high, and by the LO distribution network if it is low.
Figure 6: Power consumption of the TX, RX, and LO distribution network versus the number of MIMO radar elements. The LO power consumption is shown for two values of the power gain efficiency (PGE) $\alpha$. The power consumption of the radar front-end is dominated by the LO network for small arrays, while for large arrays it is dominated by the TX when $\alpha$ is high, and by the LO network if $\alpha$ is low.
IV MIMO Radar Back-End
IV-A MUSIC and MVDR DOA Estimation Algorithms
@@
The covariance matrix and the array matrix are used to construct a spatial spectrum function $P(\theta,\phi)$. In the MUSIC algorithm, $P(\theta,\phi)$ is related to the signal and noise subspace of the covariance matrix [25], while in the MVDR algorithm it is derived from the inverse of the covariance matrix [26]. Finally, a search algorithm is deployed to find peaks of the spatial spectrum function as DOA of the source signals.
IV-B DOA Estimation Results
@@
The MUSIC and MVDR algorithms are evaluated for various architectures of the 2D MIMO radar. The reference distance is assumed as $d_{0}={\rm{1\,m}}$. Root-mean-square error (RMSE) is used as the accuracy metric
${\rm{RMSE}}=\sqrt{\frac{1}{KM}\sum_{j=1}^{M}\sum_{i=1}^{K}\left((\hat{\phi}_{i,j}-\phi_{i})^{2}+(\hat{\theta}_{i,j}-\theta_{i})^{2}\right)},$ (31)
@@
where $(\phi_{i},\theta_{i})$ is the actual DOA of the i-th target, $(\hat{\phi}_{i,j},\hat{\theta}_{i,j})$ is the estimated DOA of the i-th target in the j-th sample of the Monte Carlo runs, $K$ is the number of targets, and $M$ is the number of the Monte Carlo runs. The noise present in the received signal is modeled as a random Gaussian variable with zero mean and a variance determined based on the SNR, i.e., $n\sim\mathcal{N}(0,\sigma^{2})$. We evaluated the simulation results for different numbers of runs and selected $M=100$ to achieve high accuracy with practical simulation times. In this work, it is assumed that there are three targets, $K=3$, with equal signal amplitudes.
IV-B1 64-Element Array
@@
For a 64-element array realized as an $8\times 8$ square array, the RMSE in terms of SNR and distance is shown in Fig. 7. The SNR and the distance are, respectively, varied in the range of 0–30 dB and 1–10 m, as typical values in indoor wireless sensing scenarios. The MUSIC algorithm outperforms the MVDR algorithm in the all SNR levels and distances, especially at long-distance and low-SNR conditions. Fig. 7(a) indicates that at short distances, e.g., 1 m, ${\rm RMSE}<1^{\circ}$ can be maintained even at very low SNR levels. This is the result of an improved SNR of the MIMO array that for the 64-element array is derived using (6) as $\rm 10\log_{10}(64)\approx 18\,dB$. As shown in Fig. 7(b), for the SNR of 20 dB, ${\rm RMSE}<0.6^{\circ}$ can be achieved across the 1–10 m distances using the MUSIC algorithm.
These results highlight that with a proper algorithm, the required SNR to achieve a specific DOA accuracy (e.g., ${\rm RMSE}<1^{\circ}$) can be greatly reduced. This leads to much lower required TX output power and, as a result, saves power consumption. The scaling of the TX output power and the RX SNR with distance provide a balance in resource allocation to the short and long distances, and further lowers the power consumption.
@@
Figure 7: RMSE for the MUSIC and MVDR algorithms with $8\times 8$ MIMO radar. For this large array, MUSIC outperforms MVSR especially in the low-SNR conditions.
IV-B2 16-Element Array
@@
For a 16-element array realized as a $4\times 4$ structure, the RMSE in terms of the SNR and distance is shown in Fig. 8. Interestingly, the MUSIC and MVDR algorithms provide very close accuracy across the ranges of SNR and distance considered in the simulations. This is in contrary with the $8\times 8$ array in which the MUSIC led to higher accuracy compared to the MVDR. The SNR improvement by the MIMO array is $\rm 10\log_{10}(16)\approx 12\,dB$. Fig. 8(b) indicates that with the 20 dB SNR, ${\rm RMSE}<2^{\circ}$ can be achieved in the 1–10 m range.
@@
Figure 8: RMSE for the MUSIC and MVDR algorithms with $4\times 4$ MIMO radar. In this case, MUSIC and MVDR provide very close results.
IV-B3 Accuracy of Adaptive Radar
In Fig. 9, the RMSE is shown versus the number of array elements. The minimum RMSE achieved using the MUSIC and MVDR algorithms, $\rm{RMSE}=\min(RMSE_{MUSIC},RMSE_{MVDR})$, is presented. It is noted that a lower RMSE can be achieved for higher SNR and shorter target distance. Increasing the number of array elements generally improves the accuracy, while the amount of improvement varies with the number of array elements, SNR, and distance. The adaptive radar architecture enables trading the accuracy with the array size to improve energy efficiency of the system. For example, when the target moves closer to the radar sensor, the desired accuracy can still be achieved using an smaller array and it is imperative to reconfigure the array structure and reduce power consumption.